# Mechanistic interpretability of large language models with applications to the financial services industry

### v1

@@ 1.2. Large Language Models (LLMs)
Recent transformer-based LLMs are systems pre-trained on an enormous amount of data with an astronomical numbers of parameters. The table below, adapted from survey [15], summarizes certain aspects of famous LLMs as of 2024, e.g. their type, number of parameters, number of tokens etc. We refer the interested reader to lecture notes [5] for an introduction to LLMs. The exposition is aimed at mathematicians and physicists, and discusses the historical pretext and the phenomenology of language models among other topics.
The advent of LLMs has stimulated conversations and posed urgent questions about this disruptive technology, including on the emergent properties of LLMs, on their alignment with human values, on the bias present in their training data, on the hallucination problem, and finally, on the challenge of interpreting LLMs. As a matter of fact, LLMs are expected to produce a rapidly growing array of risks which makes research on safety and governance mechanisms for them even more crucial [2]. In this paper, we focus on the interpretability question, beginning with a short introduction to the field of mechanistic interpretability in the next section.
@@ 2. Mechanistic interpretability
The nascent field of mechanistic interpretability seeks to “reverse engineer” neural networks. This endeavor can be construed in analogy with understanding the compiled binary program run on a virtual machine where the binary code and virtual machine/interpreter correspond to the parameters and the architecture of a neural network. In this setting, variables/memory locations roughly correspond to neurons or other “independent units” a neural network representation can be decomposed into [20]. For basic resources on the topic, we refer the reader to the guide [17], the glossary [16], the TransformerLens library [18], or the course [13].
@@ 2.1. Motivation
Despite their immense success in various tasks, the lack of transparency of LLMs presents challenges such as hallucination, toxicity, unfairness, and misalignment with human values which can hinder safe deployment of these models. Thus, there is an urgent need for a deeper understanding of the inner functioning of LLMs. Mechanistic interpretability is an important explanation technique used to this end. In this paradigm, one strives to understand an LLM at the level of neurons, circuits, and attention heads, i.e. at a micro scale (as opposed to representation engineering where the explanation occurs at a macro scale). In XAI (eXplainable AI) terminology, mechanistic interpretability can be described as a global, post-hoc, model-specific and white-box (i.e. needs access to model’s internals) approach [39]. Apart from helping to address societal risks, insights from mechanistic interpretability can help with “editing” an LLM for better performance. It can furthermore be utilized to explain training phenomena such as grokking and memorization; cf. [19].
@@ 2.2. Circuits
In analogy to cellular biology, researchers have tried to understand complicated neural networks by “zooming in” and investigating “the fundamental units” building a network, and the connections between them [21]. These building blocks are called “features”. Each feature corresponds to a “direction”, meaning a vector in the representation of a layer of the neural net. This can be an individual neuron or a linear combination of neurons in a layer. Features are connected to each other by weights to form “circuits”. More precisely, a circuit is a computational subgraph of the neural network where each edge connects two neurons/directions in adjacent layers, and comes with weights which are the weights shared between them in the network. It is claimed (admittedly speculatively) in [21] that features are typically meaningful, i.e. they correspond to articulable properties of the input; and the circuits correspond to meaningful algorithms encoded in neural networks’ weights. These ideas are perhaps better understood in the context of vision models—neurons may pick certain aspects of the input image and then group together to form more sophisticated detectors; see Figure 2.
@@
We now delve into a more precise treatment of circuits following [6, 13]. In transformers, a layer’s output is added to the residual stream, i.e. the sum of outputs of the previous layers and the original embedding. The residual stream can be considered as a “communication channel” between different transformer components such as MLP layers and attention heads; see Figure 4. The function of attention heads can then be understood as moving the information: reading from the residual stream of one token, and writing to the residual stream of another token. Following [13, chap. 1.1 §2, chap. 1.2 §4], and [6], we argue that each attention head consists of two circuits:
  • A QK (query-key) circuit “determines where to move information to and from” [13, chap. 1.1 §2]. For attention head $h$, this can be captured by the matrix product $W_{E}^{\text{\tiny T}}(W^{h}_{Q})^{\text{\tiny T}}W^{h}_{K}W_{E}$.
  • An OV (output-value) circuit “determines what information to move” [13, chap. 1.1 §2]. For attention head $h$, this can be captured by the matrix product $W_{U}W^{h}_{O}W^{h}_{V}W_{E}$.
In what appeared above, matrices $W_{E}$ and $W_{U}$ capture embedding and unembedding of tokens; $W^{h}_{K}$, $W^{h}_{Q}$ and $W^{h}_{V}$ denote the key, query and value matrices of the attention head $h$; and $W^{h}_{O}$ captures the attention head’s output weights. A thorough study of these circuits for transformers with no more than two layers and no MLP layer can be found in [6]. Great insights can be gained from investigating these toy models: when there is one transformer layer, or none, there are concrete descriptions in terms of n-gram models, whereas two-layer transformers are much more capable since now attention heads can compose (in three different ways corresponding to keys, queries, and values), and this results in creation of induction heads. Going beyond toy examples, paper [22] studies induction heads for larger models: they are implemented by circuits consisting of attention heads in different layers that work together to copy or complete patterns, and thus, contributing to in-context learning.
@@
Figure 6. In activation patching, a clean run (e.g. ”When Mary and John went to the store, John gave a drink to …” in the IOI task) is considered along with a corrupted run for which the answer is flipped (e.g. ”When Mary and John went to the store, Mary gave a drink to …”). The clean run is captured on the left. On the right, an activation from the corrupted run is patched in the corresponding one from the clean run. (Pictures courtesy of [13].)
@@ 2.3. Circuit discovery
There is a growing body of literature on circuit discovery for transformers; cf. [4]. As a concrete example, the phenomenon of grokking, i.e. sudden transition from overfitting to generalization after numerous training steps, can be explained through circuit formation. This is the content of paper [19] where grokking is reverse engineered for transformer models trained for simple arithmetic tasks. Also see [8] where grokking and “emergent” abilities in LLMs are studied through the lens of circuits.
For our purposes, we shall follow the approach of [36] which is based on algorithmic tasks, an idea that we believe carries over easily to financial services applications. The paper focuses on the Indirect Object Identification (IOI) language task with the GPT-2 Small model ([26]), a language model comprised of 12 layers and about 80 million parameters. To give an example of this task, a sentence such as ”When Mary and John went to the store, John gave a drink to …” should be completed with the word ”Mary” instead of ”John”. In that paper, a circuit (here meaning a computational subgraph) of GPT-2 Small, consisting of 26 attention heads, is identified as being responsible for the IOI task. These are further broken down into 7 categories each responsible for a human-understandable purpose, e.g., “duplicate token heads” that identify tokens already appeared in the sentence.
Algorithmic examples of the IOI instances are constructed as follows (see [36, Appendix E] and [37, IOI Dataset]):
@@
After this high-level overview of [36], we delve into the steps and tools involved in the circuit discovery carried out therein following the exposition in [13, chap. 1.3].
  • Recall that the probabilities that a language model outputs for predicting the next token are obtained from applying a softmax function to tokens’ logit values. The difference of logit values, e.g. the expression $logit$("Mary") -- $logit$("John") in the case of the IOI example above, can be utilized as a performance metric.
  • The contribution of a layer (a transformer block), and the attention heads inside it, to the residual stream can be quantified by considering the logit difference after that layer as if the subsequent layers are ignored. This direct logit attribution technique identifies heads that contribute the most to the residual stream, and can be conducted with the TransformerLens library [18]. See [13, chap. 1.3 §2] for more details.
@@
      Let us begin with the activation patching tool, the idea that was first introduced in [14]. Two different runs of the model are considered, a “clean” run and a “corrupted” one. The output answers for them are correct and incorrect respectively. We intervene on a specific activation from the corrupted run by replacing it with the corresponding activation from the clean run, and then measure the change of the output towards the correct answer. This is called denoising. Alternatively, in noising, one patches from the corrupted run into the clean run; cf. Figure 6. Through trying many different activations and assessing how much they affect the corrupted run, one can identify the activations that really matter. Finally, we point out that patching into a transformer can be done in a number of different ways, for example, into MLP layers, attention heads, or the values of the residual stream. We refer the reader to [13, chap. 1.3 §3] for more details.
@@
      The activation patching approach considers the alternative of swapping the contribution of an attention head to the residual stream with what it would have been under a different distribution while everything else remains the same. On the other hand, in path patching, one contemplates the alternative of replacing the input to an attention head from another attention head with what it would have been under a different distribution; see Figure 7. The path-patching tool, thus, investigates the importance of particular paths between a model’s components (e.g. a circuit formed by two attention heads) rather than individual model components (e.g. a single attention head). Again, a clean run and a corrupted run are involved. Looking at the IOI task again, instead of replacing ”When Mary and John went to the store, John gave a drink to …” with ”When Mary and John went to the store, Mary gave a drink to …”, three random names are used as in ”When [X] and [Y] went to the store, [Z] gave a drink to …”—hence the information about duplicate tokens is erased. We refer the reader to [13, chap. 1.3 §4] for more details.
@@ 3.1. LLM interpretability for financial services
In the financial domain, BloombergGPT is an LLM with 50 billion parameters trained on a wide range of financial data [38]. BloombergGPT is not open source, and thus, investigating it through the lens of mechanistic interpretability, e.g., circuit discovery, is not practical. Moreover, frequently retraining an LLM model like BloombergGPT is costly. Therefore, there is need for lightweight adaptations. One remarkable adaptation is the FinGPT project which facilitates prompt engineering and fine-tuning of open-source LLMs on financial data [12] for a variety of financial applications such as risk management, portfolio optimization, credit scoring, and fraud detection.
@@
  • An existing large language model can be used (through fine tuning for example) to classify customer-agent transcript with business-related tags such as “website/mobile”, “login issues”, “hard inquiry", “balance too high” etc.
@@
  • UDAAP rules) UDAAP protects consumers against the risks of harm from unfair, deceptive, or abusive practices by financial institutions. Harm does not have to be monetary and can result from increased difficulty of consumer understanding of the overall costs or risks of the product and the potential harm to the consumer associated with the product. Financial institutions monitor their customer-agent interactions. When the consumer suggests potential UDAAP (Unfair, Deceptive, or Abusive Acts or Practices) concerns, such interactions need to be identified for further review. One may use the following criteria to define UDAAP as an algorithmic financial task:
    • UDAAP Keywords: Unfair/not fair, trick/tricky/tricked, cheat/cheated/cheating, bait/switch, deceptive/deceitful, abuse/abused, mislead/misleading/misled, lie/lied/lying, fraud/defraud, taking/took advantage of me because I’m a [member of a vulnerable population] (e.g. student, uneducated).
    We design the following algorithmic examples for UDAAP circuit discovery:
@@
      DECEPTIVE-PRACTICES: The customer is interested in a personal loan and the agent mentions that they offer personal loans with APR as low as $4.99\%$. Example: ”You are [UDAAP-BEHAVIOR]. You did not disclose that the rate $4.99\%$ only applies to applicants with excellent credit scores and my rate is significantly higher. Your loans also comes with hidden origination fees that the agent did not mention."
@@
        States may have broader protected classes under their respective fair lending laws, including but not limited to military status, ancestry, or political affiliation.
@@
        ”Why are you quoting me a higher rate than my neighbor when we have the same credit score and income. It’s because I’m old!”
        ”You are denying all my requests for a payment plan because I’m a single mother.”
        ”I don’t understand why I have all these late fees. I think you’re trying to take advantage of me because I’m disabled.”
        ”I asked for a credit line increase multiple times, and you said no because I’m on unemployment.”
    We design the following algorithmic examples for FL circuit discovery:
@@
      ”I don’t understand why I have these late fees. I think you’re trying to [UDAAP-VERB] me because I’m a [MEMBER-of-PROTECTED-CLASS].”[UDAAP-VERB] is given by: trick, cheat, deceit, abuse, mislead, defraud, taking advantage of.
@@
      Note that [MEMBER-of-PROTECTED-CLASS] can be an item from old person, elderly, single mother, disabled, minority, etc.
@@
        ”You are charging me a higher rate than my neighbor. It’s because I’m a [MEMBER-of-PROTECTED-CLASS];”
@@ 3.2. Numerical Experiments
Figure 9, illustrates the calculated logit difference by decomposing the residual stream after each layer (see logit lens from [18]). [^3] Recall from §2.3 that, one can interpret the result of residual stream decomposition and calculating logit differences as simulating what would happen if subsequent layers are removed after each layer [13]. As one can see the logit difference generally keeps improving through the layers. This is in contrast to the IOI task where one can localize the task completion to merely a few final layers, i.e., layers $7$, $8$, and $9$ [13]. We believe that this is due to the fact that our financial task at hand is far more complex when compared to the IOI task which involves moving information around from the indirect object and not the subject, requiring less information processing.
@@
Next, we further decompose the output of each attention layer into the sum of the outputs of the corresponding attention heads (see Figure 11). As expected, there are quite a few attention heads that play a substantial role in the compliance financial task completion. Of note, heads [^4] Note we use the convention that l.h specifies the head number h in the layer number l. There are $144(12\times 12)$ heads in total in GPT-2 Small. 11.4, 8.9, and 6.3 contribute a lot positively, while heads 0.6, 11.0, and 10.7 contribute a lot negatively to the task completion (see Figure 12).

### v2

@@ 1.2. Large Language Models (LLMs)
Recent transformer-based LLMs are systems pre-trained on an enormous amount of data with an astronomical numbers of parameters. The table below, adapted from survey [16], summarizes certain aspects of famous LLMs as of 2024, e.g., their type, number of parameters, number of tokens etc. We refer the interested reader to lecture notes [5] for an introduction to LLMs. The exposition is aimed at mathematicians and physicists, and discusses the historical pretext and the phenomenology of language models among other topics.
The advent of LLMs has stimulated conversations and posed urgent questions about this disruptive technology, including on the emergent properties of LLMs [41], on their alignment with human values [35], on the bias present in their training data [10], on the hallucination problem [43], and finally, on the challenge of interpreting LLMs [36]. As a matter of fact, LLMs are expected to produce a rapidly growing array of risks which makes research on safety and governance mechanisms for them even more crucial [2]. In this paper, we focus on the interpretability question, beginning with a short introduction to the field of mechanistic interpretability in the next section.
@@ 2. Mechanistic interpretability
The nascent field of mechanistic interpretability seeks to “reverse engineer” neural networks. This endeavor can be construed in parallel with understanding the compiled binary program run on a virtual machine where the binary code and virtual machine/interpreter correspond to the parameters and the architecture of a neural network. In this setting, variables/memory locations roughly correspond to neurons or other “independent units” a neural network representation can be decomposed into [21]. For basic resources on the topic, we refer the reader to the guide [18], the glossary [17], the TransformerLens library [19], or the course [14].
@@ 2.1. Motivation
Despite their immense success in various tasks, the lack of transparency of LLMs presents challenges such as hallucination, toxicity, unfairness, and misalignment with human values which can hinder safe deployment of these models. Thus, there is an urgent need for a deeper understanding of the inner functioning of LLMs. Mechanistic interpretability is an important explanation technique used to this end. In this paradigm, one strives to understand an LLM at the level of neurons, circuits, and attention heads, i.e., at a micro scale (as opposed to representation engineering [44, 45] where the explanation occurs at a macro scale). In XAI (eXplainable AI) terminology, mechanistic interpretability can be described as a global, post-hoc, model-specific and white-box (i.e., needs access to model’s internals) approach [44]. Apart from helping to address societal risks, insights from mechanistic interpretability can help with “editing” an LLM for better performance. It can furthermore be utilized to explain training phenomena such as grokking and memorization; cf. [20].
@@ 2.2. Circuits
By analogy with cellular biology, researchers have tried to understand complicated neural networks by “zooming in” and investigating “the fundamental units” building a network, and the connections between them [22]. These building blocks are called “features”. Each feature corresponds to a “direction”, meaning a vector in the representation of a layer of the neural net. This can be an individual neuron or a linear combination of neurons in a layer. Features are connected to each other by weights to form “circuits”. More precisely, a circuit is a computational subgraph of the neural network where each edge connects two neurons/directions in adjacent layers, and comes with weights which are the weights shared between them in the network. It is claimed (admittedly speculatively) in [22] that features are typically meaningful, i.e., they correspond to articulable properties of the input; and the circuits correspond to meaningful algorithms encoded in neural networks’ weights. These ideas are perhaps better understood in the context of vision models—neurons may pick certain aspects of the input image and then group together to form more sophisticated detectors; see Figure 2.
@@
We now delve into a more precise treatment of circuits following [6, 14]. In transformers, a layer’s output is added to the residual stream, i.e., the sum of outputs of the previous layers and the original embedding. The residual stream can be considered as a “communication channel” between different transformer components such as MLP layers and attention heads; see Figure 4. The function of attention heads can then be understood as moving the information: reading from the residual stream of one token, and writing to the residual stream of another token. Following [14, chap. 1.1 §2, chap. 1.2 §4], and [6], we argue that each attention head consists of two circuits (Figure 5 (top)):
  • A QK (query-key) circuit “determines where to move information to and from” [14, chap. 1.1 §2]. For attention head $h$, this can be captured by the matrix product $W_{E}^{\text{\tiny T}}(W^{h}_{Q})^{\text{\tiny T}}W^{h}_{K}W_{E}$.
  • An OV (output-value) circuit “determines what information to move” [14, chap. 1.1 §2]. For attention head $h$, this can be captured by the matrix product $W_{U}W^{h}_{O}W^{h}_{V}W_{E}$.
In what appeared above, matrices $W_{E}$ and $W_{U}$ capture embedding and unembedding of tokens; $W^{h}_{K}$, $W^{h}_{Q}$ and $W^{h}_{V}$ denote the key, query and value matrices of the attention head $h$; and $W^{h}_{O}$ captures the attention head’s output weights. A thorough study of these circuits for transformers with no more than two layers and no MLP layer can be found in [6]. Great insights can be gained from investigating these toy models: when there is one transformer layer, or none, there are concrete descriptions in terms of n-gram models, whereas two-layer transformers are much more capable since now attention heads can compose (in three different ways corresponding to keys, queries, and values), and this results in creation of induction heads (Figure 5 (bottom)). Going beyond toy examples, paper [23] studies induction heads for larger models: they are implemented by circuits consisting of attention heads in different layers that work together to copy or complete patterns, and thus, contributing to in-context learning.
@@
Figure 6. In activation patching, a clean run (e.g., ”When Mary and John went to the store, John gave a drink to …” in the IOI task) is considered along with a corrupted run for which the answer is flipped (e.g., ”When Mary and John went to the store, Mary gave a drink to …”). The clean run is captured on the left. On the right, an activation from the corrupted run is patched in the corresponding one from the clean run. (Pictures courtesy of [14].)
@@ 2.3. Circuit discovery
There is a growing body of literature on circuit discovery for transformers; cf. [4]. As a concrete example, the phenomenon of grokking, i.e., sudden transition from overfitting to generalization after numerous training steps, can be explained through circuit formation. This is the content of paper [20] where grokking is reverse engineered for transformer models trained for simple arithmetic tasks. Also, see [8] where grokking and “emergent” abilities in LLMs are studied through the lens of circuits.
For our purposes, we shall follow the approach of [39] which is based on algorithmic tasks, an idea that we believe carries over easily to financial services applications. The paper focuses on the Indirect Object Identification (IOI) language task with the GPT-2 Small model ([27]), a language model comprised of 12 layers and about 80 million parameters. To give an example of this task, a sentence such as ”When Mary and John went to the store, John gave a drink to …” should be completed with the word ”Mary” instead of ”John”. In that paper, a circuit (here, meaning a computational subgraph) of GPT-2 Small, consisting of 26 attention heads, is identified as being responsible for the IOI task. These are further broken down into 7 categories each responsible for a human-understandable purpose, e.g., “duplicate token heads” that identify tokens already appeared in the sentence.
Algorithmic examples of the IOI instances are constructed as follows (see [39, Appendix E] and [40, IOI Dataset]):
@@
After this high-level overview of [39], we delve into the steps and tools involved in the circuit discovery carried out therein following the exposition in [14, chap. 1.3].
  • Recall that the probabilities that a language model outputs for predicting the next token are obtained from applying a softmax function to tokens’ logit values. The difference of logit values, e.g., the expression $logit$("Mary") -- $logit$("John") in the case of the IOI example above, can be utilized as a performance metric.
  • The contribution of a layer (a transformer block), and the attention heads inside it, to the residual stream can be quantified by considering the logit difference after that layer as if the subsequent layers are ignored. This direct logit attribution technique identifies heads that contribute the most to the residual stream, and can be conducted with the TransformerLens library [19]. See [14, chap. 1.3 §2] for more details.
@@
      Let us begin with the activation patching tool, the idea that was first introduced in [15]. Two different runs of the model are considered, a “clean” run and a “corrupted” one. The output answers for them are correct and incorrect, respectively. We intervene on a specific activation from the corrupted run by replacing it with the corresponding activation from the clean run, and then measure the change of the output towards the correct answer. This is called denoising. Alternatively, in noising, one patches from the corrupted run into the clean run; cf. Figure 6. Through trying many different activations and assessing how much they affect the corrupted run, one can identify the activations that really matter. Finally, we point out that patching into a transformer can be done in a number of different ways, for example, into MLP layers, attention heads, or the values of the residual stream. We refer the reader to [14, chap. 1.3 §3] for more details.
@@
      The activation patching approach considers the alternative of swapping the contribution of an attention head to the residual stream with what it would have been under a different distribution while everything else remains the same. On the other hand, in path patching, one contemplates the alternative of replacing the input to an attention head from another attention head with what it would have been under a different distribution; see Figure 7. The path-patching tool, thus, investigates the importance of particular paths between a model’s components (e.g., a circuit formed by two attention heads) rather than individual model components (e.g., a single attention head). Again, a clean run and a corrupted run are involved. Looking at the IOI task again, instead of replacing ”When Mary and John went to the store, John gave a drink to …” with ”When Mary and John went to the store, Mary gave a drink to …”, three random names are used as in ”When [X] and [Y] went to the store, [Z] gave a drink to …”—hence the information about duplicate tokens is erased. We refer the reader to [14, chap. 1.3 §4] for more details.
@@ 3.1. LLM interpretability for financial services
In the financial domain, BloombergGPT is an LLM with 50 billion parameters trained on a wide range of financial data [42]. BloombergGPT is not open source, and thus, investigating it through the lens of mechanistic interpretability, e.g., circuit discovery, is not practical. Moreover, frequently retraining an LLM model like BloombergGPT is costly. Therefore, there is a need for lightweight adaptations. One remarkable adaptation is the FinGPT project which facilitates prompt engineering and fine-tuning of open-source LLMs on financial data [13] for a variety of financial applications such as risk management, portfolio optimization, credit scoring, and fraud detection.
@@
  • An existing large language model can be used (through fine tuning, for example) to classify customer-agent transcripts with business-related tags such as “website/mobile”, “login issues”, “hard inquiry", “balance too high” etc.
@@
  • UDAAP rules) UDAAP protects consumers against the risks of harm from unfair, deceptive, or abusive practices by financial institutions. Harm does not have to be monetary and can result from increased difficulty of consumer understanding of the overall costs or risks of the product and the potential harm to the consumer associated with the product. Financial institutions monitor their customer-agent interactions. When the consumer suggests potential UDAAP (Unfair, Deceptive, or Abusive Acts or Practices) concerns, such interactions need to be identified for further review by a financial institution. One may use the following criteria to define UDAAP as an algorithmic financial task:
    • UDAAP Keywords: Unfair/not fair, deceptive/deceitful, abusive/abuse/abused, tricky/trick/tricked, cheating/cheat/cheated, misleading/mislead/misled, lying/lie/lied, taking/took advantage of me because I’m a [member of a vulnerable population] (e.g., students and minority groups).
    Therefore, we design the following algorithmic examples for UDAAP circuit discovery:
@@
      DECEPTIVE-PRACTICES: The customer is interested in a personal loan and the agent mentions that they offer personal loans with APR as low as $4.99\%$. Example: ”You are [UDAAP-BEHAVIOR]. You did not disclose that the rate $4.99\%$ only applies to applicants with excellent credit scores and my rate is significantly higher. Your loans also come with hidden origination fees that the agent did not mention."
@@
        States may have broader protected classes under their respective fair lending laws, including but not limited to military status, ancestry, domestic violence victim status, or political affiliation.
@@
        ”I requested for a credit line increase multiple times, and you said no because I’m on unemployment.”
        ”Why are you denying all my requests for a payment plan. It’s because I’m a single mother.”
        ”I don’t understand why I have all these late and hidden fees. I think you’re trying to take advantage of me because I’m disabled.”
        ”You are quoting me a much higher rate than my neighbor when we have the same credit score and income. It’s because I’m old!”
    We design the following algorithmic examples for Fair Lending circuit discovery:
@@
      ”I don’t understand why I have all these late fees. I think you’re trying to [UDAAP-VERB] me because I’m a [MEMBER-of-PROTECTED-CLASS].”[UDAAP-VERB] is given by: trick, cheat, deceit, abuse, mislead, defraud, taking advantage of.
@@
      Note that [MEMBER-of-PROTECTED-CLASS] can be an item from elderly, single mother, disabled, minority, etc.
@@
        ”You are charging me a much higher rate than my neighbor. It’s because I’m a [MEMBER-of-PROTECTED-CLASS];”
@@ 3.2. Numerical Experiments
Figure 9, illustrates the calculated logit difference by decomposing the residual stream after each layer (see logit lens from [19]). [^3] Recall from §2.3 that, one can interpret the result of residual stream decomposition and calculating logit differences as simulating what would happen if subsequent layers are removed after each layer [14]. As one can see the logit difference generally keeps improving through the layers. This is in contrast to the IOI task where one can localize the task completion to merely a few final layers, i.e., layers $7$, $8$, and $9$ [14]. We believe that this is due to the fact that our compliance task at hand is far more complex when compared to the IOI task which involves moving information around from the indirect object and not the subject, requiring less information processing.
@@
Next, we further decompose the output of each attention layer into the sum of the outputs of the corresponding attention heads (see Figure 11). As expected, there are quite a few attention heads that play a substantial role in the compliance financial task completion. Of note, heads [^4] Note we use the convention that l.h specifies the head number h in the layer number l (indexing starts from 00). There are $144(12\times 12)$ heads in total in GPT-2 Small. 11.4, 8.9, and 6.3 contribute a lot positively, while heads 0.6, 11.0, and 10.7 contribute a lot negatively to the task completion (see Figure 12).